OpenStudy (anonymous):

A toy is made from red and yellow bricks number of red bricks : number of yellow bricks = 5 : 2 there are 210 more red bricks than yellow bricks How many red bricks are in the toy?

OpenStudy (anonymous):

red=x+120 yellow=x 2(x+120)=5x 2x+240=5x x=80